OpenStudy (anonymous):

There are 10 balls, numbered 1-10, in a box. a ball is drawn and replaced. A second is drawn. what is the probability that the first is a 3 and the second is NOT 9?

OpenStudy (rulnick):

a. P(first is 3) = 1/10 b. P(second is 9) = 1/10 c. P(second is NOT 9) = 1-P(second is 9) Multiply (a) by (c) to get your answer.

OpenStudy (rulnick):

Get it?

OpenStudy (anonymous):

1/10*9/10?